OpenStudy (anonymous):

The similar triangular ramps used for Kevin’s tree jump have a ratio of 5:4. If the height of the first ramp is 20 feet, what is the height of the second ramp? If the area of the smaller ramp is 16ft2, what is the area of the larger ramp?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

5/4 = 20/x solve for x

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

x = 5

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

thats incorrect

jimthompson5910 (jim_thompson5910):

5/4 = 20/x 5*x = 4*20 5x = 80 x = ??

OpenStudy (anonymous):

llet me do this i can do it

OpenStudy (anonymous):

16?

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

that's the height of the second ramp

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

do you know how to do the second part?

OpenStudy (anonymous):

no:(

OpenStudy (anonymous):

wait do i just double it?

jimthompson5910 (jim_thompson5910):

ok you need to solve for y (5/4)^2 = y/16

OpenStudy (anonymous):

1.5625

jimthompson5910 (jim_thompson5910):

(5/4)^2 = y/16 (5^2)/(4^2) = y/16 25/16 = y/16 ... ... ... y = ??

OpenStudy (anonymous):

confused....

OpenStudy (anonymous):

it's not a rational #

jimthompson5910 (jim_thompson5910):

25/16 = y/16 25*16 = y*16 400 = 16y 16y = 400 y = ??

OpenStudy (anonymous):

oh okay hold on:)

OpenStudy (anonymous):

25

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

yes y = 25

OpenStudy (anonymous):

done?

jimthompson5910 (jim_thompson5910):

good, so the area of the larger ramp is 25 square feet

OpenStudy (anonymous):

thank you baby

jimthompson5910 (jim_thompson5910):

sure thing